zarkam21 (zarkam21):

What is 2b+U=28 in slope-intercept form?

OpenStudy (baby_bear69):

subtract 2b from each side, and you have u = -2b + 28, giving you a y-intercept of 28 and a slope of negative 2.

OpenStudy (baby_bear69):

So what would the slope intercept equation be?

zarkam21 (zarkam21):

Umm 28=-2

OpenStudy (anonymous):

In the form of y = mx + b.... ._.

zarkam21 (zarkam21):

Y=-2x+28

OpenStudy (anonymous):

Yush.

OpenStudy (baby_bear69):

My computer froze :p sorry. but yes that's right!

zarkam21 (zarkam21):

So for the equation b+u=18, the slope intercept form would be y=18x

OpenStudy (baby_bear69):

subtract b from each side of the equation, and you have u = -b + 18. This means that the y-intercept is 18, and the slope is negative 1.

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18

zarkam21 (zarkam21):

So Y=-1x+18